# Order and disorder in geometrically frustrated magnets

### v1

@@ I Introduction
In 1973, Anderson proposed [1] that the ground state of the Heisenberg model with nearest-neighbor antiferromagnetic (AF) interactions on a triangular lattice may lack magnetic order and instead display liquid-like properties with strong quantum fluctuations. Although this proposal turned out to be incorrect (the model develops the so-called $120^{\circ}$ order  [2, 3, 4]), it initiated a series of theoretical studies [5, 6, 7] of “quantum spin liquids” (QSLs), magnetic states that lack long-range order (LRO) at $T=0$ and, broadly speaking, exhibit liquid-like properties due to strong quantum fluctuations.
@@
The term “spin liquid” commonly refers to a $T=0$ state of a magnetic material that lacks LRO. The term “quantum spin liquid” is used to label a quantum-coherent state in which, in addition to the lack of LRO, the quantum nature of spins gives rise to qualitatively new properties. Despite some commonalities among the different QSL models, in our view, there is not one universal, commonly accepted definition of a QSL. Often, in theoretical work, a spin liquid is referred to as a QSL if it displays long-range entanglement or topological properties, such as topological order or fractionalized excitations, or can be described in terms of emergent gauge fields.
@@
The aim of this review is to outline a minimal set of existing facts that apply to all GF magnets in the hope of informing future theoretical work. Thus, we will restrict discussion to the most common experimental quantities, because such data allow for intercomparison among all GF materials. While GF magnets have been experimentally probed in many different ways, the most common quantities are the magnetic susceptibility [$\chi(T)$], specific heat [$C(T)$], and the intensity of neutron scattering. Other, more targeted probes include NMR, $\mu$SR, and optical techniques, which will not be considered here. For more information on all of the experimental probes of GF materials we refer the reader to the 2011 review of GF magnets [25], and to more recent reviews of QSL candidates [7, 26, 32, 33].
It is useful to remember what $\chi(T)$, $C(T)$, and neutron scattering reveal about the state of a magnetic material. The susceptibility, $\chi(T)$, is a long wavelength probe that is especially useful at high temperatures for determining the spin-spin interaction strength via Curie-Weiss analysis and at low temperatures for detecting transitions, e.g., to an LRO or a spin-glass state. Most importantly for this work, $\chi(T)$ can detect non-magnetic (singlet) excitations only weakly. The specific heat, however, is sensitive to all excitations with a spin origin, including non-magnetic excitations, but also to those not spin-derived, such as phonons and charged excitations. A powerful aspect of the $C(T)$ probe is its ability to count degrees of freedom, provided the non-spin-derived excitations are accurately determined. Neutrons are, like $\chi(T)$, sensitive mainly to magnetic excitations, but they can probe these excitations over a large range of wavelengths, from mesoscopic distances down to interatomic length scales. The complementarity of these three probes is sufficient to demonstrate key universal features of GF magnets. As we argue, among those universal features is the short-range magnetic order (SRO) that forms below a certain characteristic temperature, which is significantly exceeded by the spin-spin interaction energy.
@@ II.1 Types of quenched disorder
In order to describe GF materials in the pure limit, one should understand their response to quenched disorder. One may distinguish here between two generic types of disorder: intrinsic and extrinsic. Intrinsic disorder occurs due to an unusual feature of the structure, such as in $Y_{2}Mo_{2}O_{7}$, where Jahn-Teller distortions of the $Mo$ octahedra occur randomly, resulting in randomness in the spin-spin interaction, and a spin-glass (SG) transition at $T_{g}=22K$ [44, 45]. Extrinsic disorder refers to foreign atoms introduced in the clean material. Often, such disorder is represented by vacancies, i.e. foreign non-magnetic atoms substituting for the magnetic atoms of the clean material. Of course, non-magnetic vacancies will be at least subtly different, e.g. in atomic size, and therefore will not bond with their neighboring cations in a manner identical to the magnetic ion, hence one expects small steric distortion around the vacancies. Extrinsic disorder may also be introduced by substitution on either a non-magnetic cation site or the anion site, which can lead to charge disorder or sterically induced randomness in the exchange interactions between magnetic ions [46].
@@ II.2 Importance of spatial dimensionality
This review focuses on strongly GF systems with Heisenberg spins. Ising spins on frustrating lattices present a different and informative complement to Heisenberg spins. In early theoretical work, Wannier considered 2D Ising spins on a triangular lattice and showed that, whereas ferromagnetic spin-spin interactions lead to LRO, antiferromagnetic interactions cause a total suppression of the transition and lead to a nonzero ground-state entropy (per spin) [22]. The existence of such non-vanishing ground-state entropy for theoretical Ising models makes them ”infinitely frustrated”, a situation reflected experimentally by the lack of Ising-based magnets for which the f-ratio can be defined by a measurable low-energy scale.
@@ III.1 The Curie-Weiss law
The dependence of the inverse magnetic susceptibility $\chi(T)$ on temperature $T$ for several GF materials with large $f$-ratios $f>10$ is shown in Fig. 1. At sufficiently high temperatures, GF materials exhibit the Curie-Weiss behaviour of the susceptibility
@@
Figure 1: The dependence of the inverse susceptibility $\chi^{-1}(T)$ on the temperature $T$ for several GF materials. The susceptibility and the temperature are shown, respectively, in units of $C/\theta_{W}$ and $\theta_{W}$, where $C$ and $\theta_{W}$ are the Curie and Weiss constants. The upper and lower insets, represented by the boxes in the main frame, show, respectively, the spin-glass transitions in $NiGa_{2}S_{4}$ and $Na_{4}Ir_{3}O_{8}$ and $SrCr_{8.8}Ga_{3.2}O_{19}$.
@@ IV Hidden energy scale and short-range order
In Ref. [81], the spin-glass transition temperature $T_{g}$ has been analyzed as a function of the density of vacancy defects in all frustrated materials for which the data on such transitions is available. Figure 2 summarizes the discovered trends.
@@
In the limit of vanishing density concentration, the glass transition temperature extrapolates to a scale of order $T^{*}\sim 10K$, which we dubbed the “hidden energy scale”, in all analyzed frustrated compounds 111with the exception of the compounds $Fe_{3-3x}Ga_{3x-1}Ti_{5}$, in which a significantly higher glass-transition temperature may be attributable to the formation of antiferromagnetic clusters [94, 135]. This scale is significantly exceeded by the Weiss constant $\theta_{W}$, which characterizes the strength of spin-spin interactions and is typically of order several hundred Kelvin [81].
@@ IV.2 Neutron data
We note that for SG transitions in conventional, non-GF materials neutron scattering also shows an increase in diffuse scattering on cooling towards the glass transition temperature [62, 95]. In these systems, however, this magnetic scattering is caused by the degrees of freedom that undergo SG freezing. In GF systems, on the other hand, neutrons are scattered by the background spins. Neutron scattering cannot, of course, differentiate between the two scenarios, but in GF systems, it clearly evidences the SRO established by the bulk of the spins.
@@ IV.3 A scenario for the hidden energy scale
We note that while the two-peak feature is a universal property of GF magnets, only one of those peaks is usually observed due to an insufficient temperature range accessed in the measurements. In GF materials with 3d elements, the higher-temperature peak lies at temperatures $\theta_{W}\gtrsim 100K$ and is rarely captured in experiments. In GF materials with rare-earth elements, as discussed earlier, the higher-temperature peak may occur at significantly lower temperatures, exemplified by $YbMgGaO_{4}$ [119] and possibly $Ce_{2}Zr_{2}O_{7}$ [120] with the higher-temperature peaks located, respectively, near $2K$ and $0.2K$, and are, therefore, readily detectable. The lower-temperature peak in the latter material, however, is located at rather low temperatures $T\lesssim 50mK$, which may be inaccessible in experiment.
@@
associated with the lower peak of $C(T)$ matches the respective Ising ground-state entropy. If the lower-temperature peak is well separated from the higher-temperature one, comparing the entropy under it with the Ising ground-state entropy on the same lattice can be used, therefore, for verifying the described mechanism of the hidden energy scale.
The values of the Ising ground-state entropy are known for common 2D lattices, such as triangular [22, 121, 122]) and kagome [123, 124] lattices, based on analytical arguments. For other lattices, e.g. the pyrochlore lattice [55, 125], the value of the entropy can be found numerically. We note that magnetism in many GF systems comes from higher spins, such spins-$1$ or spins-$3/2$. Strictly speaking, the entropy under the lower peak in these systems has to be compared with the ground-state entropy of the respective higher-spin counterparts of the Ising model. Often, however, the ground-state entropy of such models will match the ground-state entropy of the conventional Ising model, corresponding to classical spins-$1/2$.
@@ IV.4.1 Temperature dependence: a test case
The Halperin-Saslow modes are linearly dispersive and, thus, can lead to the quadratic low-temperature dependence $C(T)\propto T^{2}$ in quasi-2D materials such as $NiGa_{2}S_{4}$ and other layered systems. A conclusive identification of the mechanism of this quadratic behaviour of the heat capacity requires further investigation.
@@ IV.4.2 Magnetic-field (in-)dependence of the specific heat
As the excitations giving rise to the lower-temperature peak in the heat capacity $C(T)$ are continuously connected to the ground states of the Ising model, the response of those states to a magnetic field depends on whether or not the respective ground states are magnetic.
@@ V Conclusions and further directions
The origin of the hidden energy scale and spin-glass transition. The SG freezing observed at a temperature of order $T^{*}$ in many rather clean GF materials raises the question of whether QSLs are achievable in GF systems or always get replaced by an SG state. This question, in part, motivated the current work. While in our view, there is not one precise, commonly accepted definition of a QSL, an SG state is considered to be incompatible with a QSL.
In most GF systems, the main source of quenched disorder, which is necessary for the existence of an SG state, is vacancy defects. It has remained a puzzle for a while what causes the SG transitions in the purest GF materials, with vacancy defects virtually eliminated. The existence of such GF transitions in rather clean GF systems seems to have created a school of thought that the SG transition is caused by a nearly immeasurable type of quenched disorder, and the SRO observed in, e.g. neutron scattering, is caused by this glass transition.
@@
Finally, while we have concentrated on analyzing the most easily accessed measurements, now that the basic spin dynamics framework is established, other more targeted spectroscopies, such as muon spin relaxation and nuclear magnetic resonance, can also be brought to bear on describing the ground states and excitations in GF materials.
VI Acknowledgements
We are grateful to S. Haravifard, P. Popp, M. Sedik, S. Sun, and S. Zhu for collaborations on the topics of this review. This work has been supported by the NSF grant DMR2218130.

### v2

# Short-range order and hidden energy scale in geometrically frustrated magnets

@@ I Introduction
In most magnetic materials at low temperatures, spins are arranged in a regular pattern, i.e. form long-range magnetic order. In some materials, however, it may be impossible to simultaneously minimize all magnetic interactions between spins, which may lead to competition between multiple magnetic states, and the material may possibly avoid long-range order.
In 1973, Anderson proposed [1] that the ground state of the Heisenberg model with nearest-neighbor antiferromagnetic (AF) interactions on a triangular lattice may lack magnetic order and instead display liquid-like properties with strong quantum fluctuations. Although this proposal turned out to be incorrect (the model develops the so-called $120^{\circ}$ order  [2, 3, 4]), it initiated a series of theoretical studies [5, 6, 7] of “quantum spin liquids” (QSLs), magnetic states that lack long-range order (LRO) at $T=0$ and, broadly speaking, exhibit liquid-like properties due to strong quantum fluctuations. The interest in QSLs has been driven by the possibility of numerous fundamental phenomena predicted in them, such as novel topological states, fractionalized excitations, and long-range entanglement, as well as the existence of anyon excitations in some QSLs, which can be used for robust quantum computation.
@@
The term “spin liquid” commonly refers to a $T=0$ state of a magnetic material that lacks LRO. The term “quantum spin liquid” is used to label a quantum-coherent state in which, in addition to the lack of LRO, the quantum nature of spins gives rise to qualitatively new properties. Despite some commonalities among the different QSL models, in our view, there is not a universal, commonly accepted definition of a QSL. Often, in theoretical work, a spin liquid is referred to as a QSL if it displays long-range entanglement or topological properties, such as topological order or fractionalized excitations, or can be described in terms of emergent gauge fields.
@@
The aim of this review is to outline a minimal set of existing facts that apply to all GF magnets in the hope of informing future theoretical work. Thus, we will restrict discussion to the most common experimental quantities, because such data allow for intercomparison among all GF materials. While GF magnets have been experimentally probed in many different ways, the most common quantities are the magnetic susceptibility [$\chi(T)$], specific heat [$C(T)$], and the intensity of neutron scattering.
Other probes include NMR and $\mu$SR. We will not consider these here because, in our view, each has a significant level of interpretive uncertainty. Each one gives good information on the internal field at the probe site, but the identity of the site is either unknown in the case of $\mu$SR [32, 33, 34] or not well defined, as shown by the two distinct signals in $NiGa_{2}S_{4}$, which were interpreted as evidence of stacking faults [32].
It is useful to remember what $\chi(T)$, $C(T)$, and neutron scattering reveal about the state of a magnetic material. The susceptibility, $\chi(T)$, is a long wavelength probe that is especially useful at high temperatures for determining the spin-spin interaction strength via Curie-Weiss analysis and at low temperatures for detecting transitions, e.g., to an LRO or a spin-glass state. Most importantly for this work, $\chi(T)$ can detect non-magnetic (singlet) excitations only weakly. The specific heat, however, is sensitive to all excitations with a spin origin, including non-magnetic excitations, but also to those not spin-derived, such as phonons and charged excitations. A powerful aspect of the $C(T)$ probe is its ability to count degrees of freedom, provided the non-spin-derived excitations are accurately determined. Neutrons, like $\chi(T)$, can probe magnetic and structural excitations and can access them over a large range of wavelengths, from mesoscopic distances down to interatomic length scales. The complementarity of these three probes is sufficient to demonstrate key universal features of GF magnets. As we argue, among those universal features is the short-range magnetic order (SRO) that forms below a certain characteristic temperature, which is significantly exceeded by the spin-spin interaction energy.
@@ II.1 Types of quenched disorder
In order to describe GF materials in the pure limit, one should understand their response to quenched disorder. One may distinguish here between two generic types of disorder: intrinsic and extrinsic.
Intrinsic disorder occurs without introducing quenched randomness in the synthesis process. For example, in $Y_{2}Mo_{2}O_{7}$, Jahn-Teller distortions of the $Mo$ octahedra result in randomness in the spin-spin interaction and a spin-glass (SG) transition at $T_{g}=22K$ [45, 46]. In some materials, different ions have to randomly occupy the same crystallographic site to ensure charge balance between cations and anions. For example, in $Fe_{2}TiO_{5}$, the $Fe^{3+}$ and $Ti^{4+}$ ions are randomly located on both $A$ and $B$ sites of the pseudo-Brookite structure [47]. In $NaCdCo_{2}F_{7}$, the $A$ site can be occupied by both $Na^{1+}$ and $Cd^{2+}$ [48]. Another example of such disorder comes from the inverse spinels, such as $B(AB)O_{4}$, in which site mixing occurs between the $A^{2+}$ and $B^{3+}$ ions [49].
Extrinsic disorder refers to foreign atoms introduced in the clean material. Often, such disorder is represented by vacancies, i.e. foreign non-magnetic atoms substituting for the magnetic atoms of the clean material. Of course, non-magnetic vacancies will be at least subtly different in atomic size and, therefore, will not bond with their neighboring cations in a manner identical to the magnetic ion, hence one expects small steric distortion around the vacancies. Extrinsic disorder may also be introduced by substitution on either a non-magnetic cation site or the anion site, which can lead to charge disorder or sterically induced randomness in the exchange interactions between magnetic ions [50].
@@ II.2 Importance of spatial dimensionality
This review focuses on strongly GF systems with Heisenberg spins. Ising spins on frustrating lattices present a different and informative complement to Heisenberg spins. In early theoretical work, Wannier considered 2D Ising spins on a triangular lattice and showed that, whereas ferromagnetic spin-spin interactions lead to LRO, antiferromagnetic interactions cause a total suppression of the transition and lead to a nonzero ground-state entropy (per spin) [22]. The existence of such non-vanishing ground-state entropy for theoretical Ising models makes them ”infinitely frustrated”, a situation reflected experimentally by the lack of Ising-based magnets for which the f-ratio can be defined by a measurable low-energy scale. Among the materials that have Ising spins but do not exhibit an ice-like state are $Co_{3}Mg(OH)_{6}Cl_{2}$ [58] and $A_{2}Co(SeO_{3})_{2}$, where $A=K$ or $Rb$ [59].
@@ III.1 The Curie-Weiss law
The dependence of the inverse magnetic susceptibility $\chi(T)$ on temperature $T$ for several GF materials [66, 67, 68, 69, 70, 71, 72, 73, 36] with large $f$-ratios $f>10$ is shown in Fig. 1. At sufficiently high temperatures, GF materials exhibit the Curie-Weiss behaviour of the susceptibility
@@
Figure 1: The dependence of the inverse susceptibility $\chi^{-1}(T)$ on the temperature $T$ for several GF materials [66, 67, 68, 69, 70, 71, 72, 73, 36]. The susceptibility and the temperature are shown, respectively, in units of $C/\theta_{W}$ and $\theta_{W}$, where $C$ and $\theta_{W}$ are the Curie and Weiss constants. The upper and lower insets, represented by the boxes in the main frame, show, respectively, the spin-glass transitions in $NiGa_{2}S_{4}$ and $Na_{4}Ir_{3}O_{8}$ and $SrCr_{8.8}Ga_{3.2}O_{19}$. The dashed line corresponds to the Curie-Weiss law.
@@ IV Hidden energy scale and short-range order
In Ref. [94], the spin-glass transition temperature $T_{g}$ has been analyzed as a function of the density of vacancy defects in all frustrated materials for which the data on such transitions are available. Figure 2 summarizes the discovered trends.
@@
In the limit of vanishing density concentration, the glass transition temperature extrapolates to a scale of order $T^{*}\sim 10K$, which we dubbed the “hidden energy scale”, in all analyzed frustrated compounds 111with the exception of the compounds $Fe_{3-3x}Ga_{3x-1}Ti_{5}$, in which a significantly higher glass-transition temperature may be attributable to the formation of antiferromagnetic clusters [47, 144]. This scale is significantly exceeded by the Weiss constant $\theta_{W}$, which characterizes the strength of spin-spin interactions and is typically a few hundred Kelvin [94].
@@ IV.2 Neutron data
We note that for SG transitions in conventional, non-GF materials, neutron scattering also shows an increase in diffuse scattering on cooling towards the glass transition temperature [75, 107]. In these systems, however, this magnetic scattering is caused by the degrees of freedom that undergo SG freezing. In GF systems, on the other hand, neutrons are scattered by the background spins. Neutron scattering cannot, of course, differentiate between the two scenarios, but in GF systems, it clearly evidences the SRO established by the bulk of the spins.
@@ IV.3 A scenario for the hidden energy scale
We note that while the two-peak feature is a universal property of GF magnets, only one of those peaks is usually observed due to an insufficient temperature range accessed in the measurements. In GF materials with 3d elements, the higher-temperature peak lies at temperatures $\theta_{W}\gtrsim 100K$ and is rarely captured in experiments. In GF materials with significantly lower $\theta_{W}$, such as rare-earth compounds, the lower peak may lie below the lower limit of cryogenic apparatus ($\sim 50mK$) and may thus be inaccessible in experiments.
@@
associated with the lower peak of $C(T)$ matches the respective Ising ground-state entropy for a particular lattice type. If the lower-temperature peak is well separated from the higher-temperature one, comparing the entropy under it with the Ising ground-state entropy on the same lattice can be used, therefore, for verifying the described mechanism of the hidden energy scale.
The values of the Ising ground-state entropy are known for common 2D lattices, such as triangular [22, 130, 131]) and kagome [132, 133] lattices, based on analytical arguments. For other lattices, e.g. the pyrochlore lattice [61, 134], the value of the entropy can be found numerically. We note that magnetism in many GF systems comes from higher spins, such as spins-$1$ or spins-$3/2$. Strictly speaking, the entropy under the lower peak in these systems has to be compared with the ground-state entropy of the respective higher-spin counterparts of the Ising model. Often, however, the ground-state entropy of such models will match the ground-state entropy of the conventional Ising model, corresponding to classical spins-$1/2$.
@@ IV.4.1 Temperature dependence: a test case
Because the Halperin-Saslow modes are linearly dispersive, they can lead to the quadratic low-temperature dependence $C(T)\propto T^{2}$ in quasi-2D materials such as $NiGa_{2}S_{4}$ and other layered systems. A conclusive identification of the mechanism of this quadratic behaviour of the heat capacity requires further investigation.
@@ IV.4.2 Magnetic-field (in-)dependence of the specific heat
Because the excitations giving rise to the lower-temperature peak in the heat capacity $C(T)$ are continuously connected to the ground states of the Ising model, the response of those states to a magnetic field depends on whether or not the respective ground states are magnetic.
@@ V Conclusions and further directions
The origin of the hidden energy scale and spin-glass transition. The SG freezing observed at a temperature of order $T^{*}$ in many rather clean GF materials raises the question of whether QSLs are achievable in GF systems or always get replaced by an SG state. This question, in part, motivated the current work. While in our view, there is not a precise, commonly accepted definition of a QSL, an SG state is considered to be incompatible with a QSL.
In most GF systems, the main source of quenched disorder, which is necessary for the existence of an SG state, is vacancy defects. It has remained a puzzle for a while what causes the SG transitions in the purest GF materials, with vacancy defects virtually eliminated. The existence of such SG transitions in rather clean GF systems seems to have created a school of thought that the SG transition is caused by a nearly immeasurable type of quenched disorder, and the SRO observed in, e.g. neutron scattering, is caused by this glass transition.
@@
Finally, while we have concentrated on analyzing the most easily accessed measurements, now that the basic spin dynamics framework is established, other more targeted spectroscopies, such as muon spin relaxation and nuclear magnetic resonance, can also be brought to bear on describing the ground states and excitations in GF materials. Of particular interest are the dynamics of excitations below the hidden energy scale, which both NMR and $\mu$SR have shown signs of [32, 143] and may help to inform the microscopic theory of low-energy states.
VI Data Availability
The data used to generate Figs. 1, 4 and 5 are available from the cited articles, except for the data for the magnetic susceptibility in $SrCr_{9p}Ga_{12-9p}O_{19}$ with $p=0.98$ and $p=0.67$. The latter data are available from the authors on request. Figures 2 and 3 have been adapted with modification from the cited articles.
VII Conflicts of Interest
The authors declare that they have no conflicts of interest.
VIII Acknowledgements
We are grateful to S. Haravifard, P. Popp, M. Sedik, S. Sun, and S. Zhu for collaborations on the topics of this review. We also appreciate useful discussions with R. Colman, P. Prelovs̆ek, B. Sbierski, and M. Ulaga. This work has been supported by the NSF grant DMR2218130.